# RandOhm: Mitigating Impedance Side-channel Attacks using Randomized Circuit Configurations

### v2

@@ 1. Introduction
Our Contribution. To answer the above questions, we introduce RandOhm, a new approach that utilizes an end-to-end modular MTD strategy based on the partial reconfiguration of FPGAs and programmable SoCs to mitigate impedance side-channel attacks. Compared to existing reconfiguration-based mitigation methods, RandOhm generates randomized partial bitstreams once during the design phase and deploys them during runtime using an open-source bitstream manipulation tool to expedite the process, improve memory utilization on the FPGA, and provide real-time protection. By randomizing the placement and routing of circuitry through runtime reconfiguration of secret-sensitive parts, we can decorrelate data-dependent computation from impedance values, significantly reducing the information leakage through the PDN’s impedance. To show the effectiveness of our approach, we use RandOhm on 28-nm FPGAs and SoCs to protect the AES cipher implementations. We assess the resiliency of our proposed solution by performing non-profiled and profiled impedance analysis attacks. Finally, we investigate the overhead of our mitigation in terms of delay and performance.
@@ 2.1. Impedance Side-channel Attacks
The impedance of an electrical element is a frequency-dependent complex number $Z(f)$ and is represented with real and imaginary parts or in polar form $|Z|\angle\theta$. Impedance is measured at a set of different frequencies in a wide range. A common practice is to make use of vector network analyzers (VNAs) to perform scattering parameter (S-Matrix) measurements and extract impedance values. To conduct a scattering parameter measurement, VNA is connected to the target device’s PDN and a series of frequency points are set for the measurements. During the measurement, RF sine waves with specific power are generated in those frequency points and are injected into the PDN. At the same time, the reflected (or transmitted in the case of 2-port measurements) signals are received by the VNA, and the relative amplitude and phase at each frequency point are recorded. $S_{11}$ describes the scattering reflection rate of the element. In other words, it quantifies the portion of reflected RF waves ($S_{11}=V^{-}/V^{+}$). Upon measurement of $S_{11}$, a simple transformation can be used to extract the impedance profile using $Z_{DUT}=Z_{0}(1+S_{11})/(1-S_{11})$.
@@ 3.2. Target Slice Multiplexer
In Fig. 3, with the assumption that the initial target data is securely stored and streamed into the functional block (highlighted with red), a linear feedback shift register (LFSR) serves as a random selector for existing target shift-register in different physical slices on the FPGA (denoted by blue). Once the mechanism is activated, a single shift-register is chosen to load the data where other instances are cleared simultaneously. Furthermore, it is possible to re-activate the mechanism by including a trigger signal in the design.
@@ 6.2. Profiled Attack
Here, we perform TIMA (Monfared et al., 2023) against the masked AES implementation. We initiate the attack at the first clock cycle when the shares of the first key byte and the first input byte shares are loaded into the target. This approach allows TIMA to directly attack the key (share) registers, bypassing any masked operations in subsequent clock cycles. During the profiling phase, we conduct two sets of experiments in 1GHz-3GHz frequency range with 5000 frequency points: 1) we collect 10,000 traces to create templates for the masked key registers without utilizing RandOhm. 2) Then we activate RandOhm and perform the profiling stage for 100,000 traces. TIMA profiling is conducted independently for each bit of all key shares. Note that the shares are generated in a uniformly random manner, and each trace contributes to template of all target bits. Specifically, for templating each target bit, we approximately have 10,000 traces where a target bit of the concerned share is 0b0 in the first scenario. After the profiling stage, the adversary attempts to guess the key based on a limited number of attack traces. Here, to analyze the information leakage, we use a simple Difference of Mean (DM) metric. This means that for an specific profiled target key bit we have: ($DM_{b_{t}}=Abs(Mean(Tr|b_{t}=0b0)-Mean(Tr|b_{t}=0b1))$). If $DM$ is large enough to be distinguished (in terms of relative SNR), the adversary can effectively perform the template attack.
Fig. 7 shows the the impedance DM leakage for two target bits of masked keys in frequency domain. It is clearly observed that RandOhm contributes to impedance profile randomization. Note that DM leakage is averaged over 100,000 traces for MTD-enabled implantation compared to 10,000 traces in the regular AES-DOM implementation.
@@ Acknowledgment
This effort was sponsored by NSF under the grant number 2338069.

### v3

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: IEEE/ACM International Conference on Computer-Aided Design; October 27–31, 2024; New York, NY, USAbooktitle: IEEE/ACM International Conference on Computer-Aided Design (ICCAD ’24), October 27–31, 2024, New York, NY, USAdoi: 10.1145/3676536.3676687isbn: 979-8-4007-1077-3/24/10
@@ 1. Introduction
Our Contribution. To answer the above questions, we introduce RandOhm, a new approach that utilizes an end-to-end modular MTD strategy based on the partial reconfiguration of FPGAs and programmable SoCs to mitigate impedance side-channel attacks. Compared to existing reconfiguration-based mitigation methods, RandOhm generates randomized partial bitstreams once during the design phase and deploys them during runtime using an open-source bitstream manipulation tool to expedite the process, improve memory utilization on the FPGA, and provide real-time protection. By randomizing the placement and routing of circuitry through runtime reconfiguration of secret-sensitive parts, we can decorrelate data-dependent computation from impedance values, significantly reducing the information leakage through the PDN’s impedance. To show the effectiveness of our approach, we use RandOhm on 28-nm FPGAs and SoCs to protect the AES cipher implementations. We assess the resiliency of our proposed solution by performing non-profiled and profiled impedance analysis. Finally, we investigate the overhead of our mitigation in terms of delay and performance.
Source Code Availability. We publish the source code of RandOhm in : https://github.com/vernamlab/RandOhm
@@ 2.1. Impedance Side-channel Attacks
The impedance of an electrical element is a frequency-dependent complex number $Z(f)$ and is represented with real and imaginary parts or in polar form $|Z|\angle\theta$. Impedance is measured at a set of different frequencies in a wide range. A common practice is to make use of vector network analyzers (VNAs) to perform scattering parameter (S-Matrix) measurements and extract impedance values. To conduct a scattering parameter measurement, VNA is connected to the target device’s PDN and a series of frequency points are set for the measurements. During the measurement, RF sine waves with specific power are generated in those frequency points and are injected into the PDN. At the same time, the reflected signals are received by the VNA, and the relative amplitude and phase at each frequency point are recorded. $S_{11}$ describes the scattering reflection rate of the element. In other words, it quantifies the portion of reflected RF waves ($S_{11}=V^{-}/V^{+}$). Upon measurement of $S_{11}$, a simple transformation can be used to extract the impedance profile using $Z_{DUT}=Z_{0}(1+S_{11})/(1-S_{11})$.
@@ 3.2. Target Slice Multiplexer
In Fig. 3, with the assumption that the initial target data is securely stored and streamed into the functional block (highlighted with red), a linear feedback shift register (LFSR) is securely initialized with a TRNG and serves as a random selector for existing target shift-register in different physical slices on the FPGA (denoted by blue). Once the mechanism is activated, a single shift-register is chosen to load the data where other instances are cleared simultaneously. Furthermore, it is possible to re-activate the mechanism by including a trigger signal in the design.
@@ 6.2. Profiled Attack
Here, we perform TIMA (Monfared et al., 2023) against the masked AES implementation. We initiate the attack at the first clock cycle when the shares of the first key byte and the first input byte shares are loaded into the target. This approach allows TIMA to directly attack the key (share) registers, bypassing any masked operations in subsequent clock cycles. During the profiling phase, we conduct two sets of experiments in 1GHz-3GHz frequency range with 5000 frequency points: 1) we collect 10,000 traces to create templates for the masked key registers without utilizing RandOhm. 2) Then we activate RandOhm and perform the profiling stage for 100,000 traces. TIMA profiling is conducted independently for each bit of all key shares. Note that the shares are generated in a uniformly random manner, and each trace contributes to template of all target bits. Specifically, for templating each target bit, we approximately have 10,000 traces where a target bit of the concerned share is 0b0 in the first scenario. After the profiling stage, the adversary attempts to guess the key based on a limited number of attack traces. Here, to analyze the information leakage, we use a simple Difference of Mean (DM) metric. This means that for a specific profiled target key bit we have: ($DM_{b_{t}}=Abs(Mean(Tr|b_{t}=0b0)-Mean(Tr|b_{t}=0b1))$). If $DM$ is large enough to be distinguished (in terms of relative SNR), the adversary can effectively perform the template attack.
Fig. 7 shows the the impedance DM leakage for two target bits of masked keys in frequency domain. It is clearly observed that RandOhm contributes to impedance profile randomization. Note that DM leakage is averaged over 100,000 traces for MTD-enabled implementation compared to 10,000 traces in the regular AES-DOM implementation.
@@ Acknowledgment
This effort was sponsored by NSF Grant CNS-2338069.